OpenStudy (anonymous):

Maximize subject to P P=3x+2y 6x+3y≤24 3x+6y≤30 x,y ≥0

OpenStudy (anonymous):

6x+3y≤24 ---(1) 3x+6y≤30 ---(2) (1)+(2): 9x+9y≤54 x+y≤6 x≤6-y put it into 6x+3y≤24 6(6-y)+3y≤24 36-3y≤24 4≤y x≤6-y = 2 Maximum value of P = 3(4)+2(2) = 14

OpenStudy (anonymous):

Thank you.